OpenStudy (anonymous):

f(x)=x^2+2x-1 Find the two values of x that make f(x)=0. Round your answers to two decimal places.

OpenStudy (anonymous):

I can't factor it, I tried the complete the square approach... I can't figure this out.

OpenStudy (anonymous):

You can complete the square here: What will give you at least x^2 + 2x ...as a product of 2 binomials.

OpenStudy (anonymous):

(x+1)^2, I got that. But when I continued with it. I had to find the square root of 0. which I know is not right.

OpenStudy (anonymous):

Right and that will give you x^2 + 2x + 1, in order to make it a -1, what must you add?

OpenStudy (anonymous):

use this \[\Large x = \frac{b \pm \sqrt{b^{2}-4ac}}{2a}\]

OpenStudy (anonymous):

-2.

OpenStudy (anonymous):

Correct. Now get that to the other side of the equation and you have: \[(x+1)^2 = 2\] You see?

OpenStudy (anonymous):

sorry theres - b

OpenStudy (anonymous):

hmm... for some reason I didn't think the quadratic formula would be appropriate.... not sure why...

OpenStudy (anonymous):

But where would the -2 come from??

OpenStudy (anonymous):

Because you need to have a factor that will give you your original term back. -2 + 1 = -1

OpenStudy (anonymous):

Otherwise you've completely screwed the whole equation. They are equivalent expressions.

OpenStudy (anonymous):

The quadratic formula will also work here, of course. This is just another avenue to take.

OpenStudy (anonymous):

I still don't totally grasp it, but it does sound familiar... is that something your supposed to do every time?? I thought when you completed the square you just added the "new number" to both sides of the equal sign.

OpenStudy (anonymous):

You end up with the same answer finally @a.galvan2

OpenStudy (anonymous):

Which would be easier? quadratic function? or completing the square?

OpenStudy (anonymous):

Well thats correct, yes. (x+1)^2 = x^2 + 2x + 1, but that is not identical to the original expression. We have to turn that 1 into a -1. To do that we add a -2.

OpenStudy (anonymous):

So we have: \[x^2 + 2x + 1 - 2 = 0 \rightarrow (x+1)^2 -2 = 0\]

OpenStudy (anonymous):



OpenStudy (anonymous):